OpenStudy (anonymous):

Question 1.1. What is the first step in solving x2 + 2x + 3 = 0 by completing the square? (Points : 1) Factor the left side. Add 1 to each side of the equation. Divide both sides of the equation by 2. Subtract 3 from each side of the equation.

OpenStudy (anonymous):

add 2x to x which equeals 3 times 2 (multiplied by) = 6+3= 9-0= -9

OpenStudy (turingtest):

the first step to solving an equaiton of the form\[ax^2+bx+c=0\]by completing the square is by first dividing by a, giving\[x^2+\frac bax+\frac ca=0\]then adding \(\large\left(\frac b{2a}\right)^2\) to both sides. Which of your choices does that correspond to?

OpenStudy (anonymous):

A.)?

OpenStudy (anonymous):

@TuringTest

OpenStudy (anonymous):

what do you think @ZachH54

OpenStudy (anonymous):

I'm pretty sure you have to ad the x together to reach your answer.

OpenStudy (anonymous):

So A, B, C, or D?

OpenStudy (anonymous):

I now say C

OpenStudy (anonymous):

C

OpenStudy (turingtest):

what was the first step I said, technically? divide by \(a\), and what is \(a\) in this case?

OpenStudy (turingtest):

the answer is not C, by the way

OpenStudy (campbell_st):

I would say isolate all terms in x from any constants....

OpenStudy (campbell_st):

so manage the constant 1st